OpenStudy (anonymous):

Solve for y x=2y^2-6y+1

OpenStudy (lgbasallote):

you'll need two equations to get y...unless x = 0 or you're looking for y INTERCEPT..

OpenStudy (anonymous):

The original equation was y=-2x-6x+1 and I want to find the inverse

OpenStudy (anonymous):

I am stuck at x-1/2=y^2-6y

OpenStudy (anonymous):

complte the square on the right side...

OpenStudy (anonymous):

hmm, I figured...

OpenStudy (anonymous):

\[\frac{x-1}{2}+9=y^2-6y+9\]

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

And then what do I do, am I actually going to get a number for y or...?

OpenStudy (anonymous):

Okay, thanks guys!

OpenStudy (anonymous):

well that equation can be written as \(\large \frac{x−1+18}{2}=(y−3)^2\) then take the square root of both sides... and like lgb said, you'll have two expressions, positive and negative.

OpenStudy (anonymous):

yw...:)